# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v2

@@ I Introduction
The proliferation of large-scale cloud storage systems and distributed file architectures (e.g., Windows Azure Storage, Google File System) has transformed disk failures from exceptional events into commonplace occurrences. To combat these failures, conventional replication strategies create multiple copies of data blocks across storage nodes, albeit at substantial storage costs. This limitation has driven the development of erasure coding techniques that achieve comparable reliability with significantly improved storage efficiency.
@@
$\rho_{w}\geqslant n_{F}-t(\overline{\Delta}+d-1)$ [27, Theorem IV.4]
@@ II Preliminary
  • For a set $S\subseteq[n]$, let $\mathcal{C}|_{S}$ denote the punctured code formed by deleting symbols $c_{j}$ for all $j\in[n]\setminus S$ from all codewords of $\mathcal{C}$;
@@ II-A Locally Repairable Codes
Let $\mathcal{C}$ be an $n,k,d]_{q}$ code. The $i$th code symbol $c_{i}$ of $\mathcal{C}$ is said to have $(r,\delta)$-locality if there exists a subset $S_{i}\subset[n]$ (an $(r,\delta)$-repair set) satisfying
@@
where $\lambda_{j}\in\mathbb{F}_{q}^{*}$ for all $j\in\mathrm{Loc}(i)$. Note that, for given $i\in[n]$, $\mathrm{Loc}(i)$ may be not unique. In the following, we use the notation $\mathrm{Loc}(i)$ to denote one of such sets for the $i$th code symbol $c_{i}$.
@@
Lemma 3 ([28, Remark II.4]):
When $d=r+\delta$ and $(r+\delta-1)\mid n$, the Singleton-type bound (3) can be improved as
@@
Given $n,k,r,\delta,q$, let $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$ is the largest possible minimum Hamming distance of an $n$-length LRCs with $(r,\delta)_{a}$-locality for given finite field $\mathbb{F}_{q}$. When the bound in Lemma 2 or Lemma 3 is tight, then $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ or $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$. An $n,k,d]_{q}$ LRC with $(r,\delta)_{a}$-locality is called optimal if $d=d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$. To the convenience of analysis, we assume that
@@ II-B Convertible Codes
When both the initial and final codes are MDS codes, the corresponding convertible code $\bm{\mathcal{C}}$ is called an MDS convertible code. When both the initial and final codes are LRCs, the corresponding convertible code $\bm{\mathcal{C}}$ is called an locally repairable convertible code (LRCC). In this paper, we mainly consider the case where all initial codes have the same parameters in the merge regime, i.e., code length $n_{I}$, dimension $k_{I}$ and $(r_{I},\delta_{I})$-locality, which is called a $(t,1)_{q}$ LRCC with parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F},r_{F},\delta_{F})_{a}$ where the subscript $a$ means both the initial and final codes are with all symbols locality. If some parameters $r_{I},\delta_{I}$ are unknown, then we say the above LRCC is with parameters $(n_{I},k_{I},*,*;n_{F},k_{F},r_{F},\delta_{F})_{a}$.
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\overline{\Delta}\leqslant 0\text{ or }d>n_{I}-k+1{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t(k-\lceil\frac{r\overline{\Delta}}{r+1}\rceil),&\mbox{otherwise}.\end{cases}$
@@
When $r=k$, the preceding bound is equivalent to the lower bound on the access cost for MDS convertible codes proposed by [37, Theorem 9], since the $n,k,d]_{q}$ MDS codes are optimal LRCs with $(k,d)_{a}$-locality.
@@ III Improved Lower Bound on Access Cost
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. For each $1\leqslant i\leqslant t$, define
@@
    For each $1\leqslant t\leqslant\tau$, the symbols $c_{i}$ for all $i\in B_{t}$ are in the same repair set $S_{t}$ of the final code $\mathcal{C}_{F}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant j\leqslant t-1}S_{j}\right)\cap(\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1})\right|<\delta-1,$ (13)
    if $|B_{t}|<\delta-1$.
Then, for each $1\leqslant i\leqslant t$,
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta_{i}\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ k-\Delta_{i}+|B|,&\mbox{otherwise}.\end{cases}$
@@
where the first inequality holds by $|S_{t}|\leqslant r+\delta-1$ for $t\in[\tau]$, and the second inequality comes from condition (i) that $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for $1\leqslant t_{2}<t_{1}\leqslant\tau$ which implies $|\mathcal{A}_{1}|\leqslant\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$ by $|B|\leqslant(\delta-1)\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$. Note that
@@
Claim 1: We claim that the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\ell}}$ can be recovered by the symbols $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\ell})}$ for $1\leqslant\ell\leqslant\tau$. The claim can be proved by induction as follows. Denote $\overline{\mathcal{U}}\triangleq\mathcal{U}_{2,1}\cup\cdots\cup\mathcal{U}_{t,1}$. If $\ell=1\in\mathcal{A}_{2}$, then $|S_{1}\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})|<\delta-1$ by (13). Since (9) and $S_{1}$ is a repair set in $\mathcal{C}_{F}$, then $S_{1}\subseteq\mathcal{F}_{1}=\mathcal{U}_{1,1}\cup\overline{\mathcal{U}}\cup\mathcal{W}_{1}$. Furthermore, we have
@@
$\geqslant$~ $\left|(S_{\ell+1}\setminus B_{\ell+1})\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$=$~ $\left|S_{\ell+1}\setminus B_{\ell+1}\right|$
@@
where the first inequality holds by $B_{\ell+1}\subseteq S_{\ell+1}\subseteq B\cup N\subseteq M$, the first equation comes from the condition (i) that $B_{\ell+1}\cap S_{i}=\emptyset$ for $i\in[\ell]$, and the second inequality holds by condition (ii) that $|B_{\ell+1}|\leqslant\delta-1$. Then, the symbols $\bm{v}|_{B_{\ell+1}}$ and $\bm{v}|_{S_{\ell+1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{\ell+1}}$. Therefore, the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\tau}}=\bm{v}|_{B}$ can be recovered by $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\tau})}=\bm{v}|_{M\setminus B}$ according to induction.
@@
This implies that $d\leqslant n_{I}-k+1$, which is a contradiction. Thus, when $d>n_{I}-k+1$, we always have $\Delta_{1}\leqslant 0$ and therefore, $|\mathcal{D}_{1,1}|\geqslant k$, which completes the proof. ∎
The lower bound in Lemma 6 depends critically on the size of the subset $|B|$. Below, we demonstrate that there always exists a set $B$ with large cardinality.
@@
    For each $1\leqslant t\leqslant\tau$, the symbols $c_{i}$ for all $i\in B_{t}$ are in the same repair set $S_{t}$ of the code $\mathcal{C}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant i\leqslant t-1}S_{i}\right)\cap\tilde{V}\right|<\delta-1,$ (15)
    if $|B_{t}|<\delta-1$.
@@
where T=∑_1⩽j⩽ℓ-1— ~Vij—¡δ-1 —V_i_j— and the first inequality holds by the fact that in each round the set $\tilde{V}$ reduces $|V_{i_{j}}|$ for the case $|\tilde{V}_{i_{j}}|<\delta-1$ and at most $|S_{i_{j}}|\leqslant r+\delta-1$ when $|\tilde{V}_{i_{j}}|\geqslant\delta-1$.
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$ and the optimal Hamming distance $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\phi(k,r,\delta)$. Then, the write access cost is
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$
@@
$|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|\leqslant|\mathcal{U}_{i,1}|\leqslant{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{F}-d-(t-1)k+1-\phi((t-1)k,r,\delta),}$
@@
$\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1\leqslant{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{F}-2d-(t-1)k+2-\phi((t-1)k,r,\delta)}\triangleq\Delta,$
where $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$ is defined in Lemma 6. Thus, $\Delta_{i}\leqslant\Delta$ for every $i\in[t]$. According to Lemma 7, let $\tilde{V}=\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}$ then $|\tilde{V}|=\Delta_{i}+d-1$ by $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$. Then, we can find a $B$ satisfying conditions (i), (ii) and $|B|=(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$. This is to say that
@@
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,&\mbox{otherwise},\end{cases}$
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$ (23)
@@
Suppose $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$ corresponding to Lemma 3. Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $((\ell/t+h)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$ where $\delta\geqslant 2,t\mid\ell$ and $\ell/t>h\geqslant 0$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$. Then the write access cost is at least $h(r+\delta-1)$ and the read access cost is at least thrthr.
@@
Corollary 1 extends the lower bound in Theorem 1 [27, Theorem IV.4] which is only associated with $r$, to one related to $r$ and $\delta$. For the case $\delta=2$, the new lower bound is exactly the same as the one in Theorem 1. In the case where $\delta>2$, the new lower bound on the write access cost clearly outperforms the one in Theorem 1. The expressions $\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$ are not easily reducible and LRCs achieving the minimum Hamming distance bound in Lemma 2 usually require some parameter constraints [46, 8], e.g., $r+\delta-1\mid n$, or $n=w(r+\delta-1)+m,k=ur+v$ such that $0<v<r,v+\delta-1\leqslant m<r+\delta-1$, etc. Therefore, we investigate the following specific scenarios which show the new lower bound on the read access cost surpasses the one in Theorem 1, i.e., $\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\leqslant\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$:
  1. 1.
    When $\Delta\mid(r+\delta-1)$, then
  2. 2.
    For $\bm{\mathcal{C}}$ with parameters $(n_{I},kr,*,*;n_{F},tkr,r,\delta)_{a}$ and the final code $\mathcal{C}_{F}$ as an optimal LRC satisfying $n_{F}=(tk+m)(r+\delta-1)$ where $m$ is a positive integer and $tk-m\geqslant 2$, then
The reader may refer to Figure 1 as an example.
Remark 3:
Figure 1: Compare the lower bounds on read access cost of Theorem 1  [27, Theorem IV.4] and Corollary 1. We consider only the case $\Delta>0$ and $d\leqslant n_{I}-k+1$. Set the $t=3$ initial codes with parameters $n_{I}=80,k=40]_{q}$, and the final code with parameters $tk/\alpha,tk,d=15]_{q}$ and with $(r=10,\delta)_{a}$-locality where $\alpha$ is the code rate (the X-axis in the figure) of the final code. Note that the lower bound of Theorem 1 is independent of the $\delta$, therefore, the conditions for $\delta=3,4,5$ are the same blue circle in the figure.

### v3

@@ I Introduction
The proliferation of large-scale cloud storage systems and distributed file architectures (e.g., Windows Azure Storage, Google File System) has transformed disk failures from exceptional events into commonplace occurrences. To handle these failures, conventional replication strategies create multiple copies of data blocks across storage nodes, albeit at substantial storage costs. This limitation has driven the development of erasure coding techniques that achieve comparable reliability with significantly improved storage efficiency.
@@
$\rho_{w}\geqslant n_{F}-t(\overline{\Delta}+d-1)$ [42, Theorem IV.4]
@@ II Preliminary
  • For a set $S\subseteq[n]$, let $\mathcal{C}|_{S}$ denote the punctured code formed by deleting the code symbols indexed by $n]\setminus S$ from all codewords of $\mathcal{C}$;
@@ II-A Locally Repairable Codes
Let $\mathcal{C}$ be an $n,k,d]_{q}$ code. The $i$th code symbol of $\mathcal{C}$ is said to have $(r,\delta)$-locality if there exists a subset $S_{i}\subset[n]$ (an $(r,\delta)$-repair set) satisfying
@@
where $\lambda_{j}\in\mathbb{F}_{q}^{*}$ for all $j\in\mathrm{Loc}(i)$, for each codeword $\mathbf{c}=(c_{1},c_{2},\dots,c_{n})\in\mathcal{C}$. Note that, for given $i\in[n]$, $\mathrm{Loc}(i)$ may be not unique. In the following, we use the notation $\mathrm{Loc}(i)$ to denote one of such sets for the $i$th code symbol.
@@
Lemma 3 ([48, Remark II.4]):
When $r=d-\delta$ and $(r+\delta-1)\mid n$, the Singleton-type bound (3) can be improved as
@@
Given $n,k,r,\delta,q$, let $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$ is the largest possible minimum Hamming distance of an $n,k]_{q}$ LRCs with $(r,\delta)_{a}$-locality for given finite field $\mathbb{F}_{q}$. When the bound in Lemma 2 or Lemma 3 is tight, then $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ or $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$, respectively. An $n,k,d]_{q}$ LRC with $(r,\delta)_{a}$-locality is called optimal if $d=d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)$. To the convenience of analysis, we assume that
@@ II-B Convertible Codes
When both the initial and final codes are MDS codes, the corresponding convertible code $\bm{\mathcal{C}}$ is called an MDS convertible code. When both the initial and final codes are LRCs, the corresponding convertible code $\bm{\mathcal{C}}$ is called an locally repairable convertible code (LRCC). In this paper, we mainly consider the case where all initial codes have the same parameters in the merge regime, i.e., code length $n_{I}$, dimension $k_{I}$ and $(r_{I},\delta_{I})$-locality, which is called a $(t,1)_{q}$ LRCC with parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F},r_{F},\delta_{F})_{a}$, where the subscript $a$ means both the initial and final codes are with all symbols locality. If some parameters $r_{I},\delta_{I}$ are unknown, then we say the above LRCC is with parameters $(n_{I},k_{I},*,*;n_{F},k_{F},r_{F},\delta_{F})_{a}$.
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\overline{\Delta}\leqslant 0\text{ or }d>n_{I}-k+1,\\ t(k-\lceil\frac{r\overline{\Delta}}{r+1}\rceil),&\mbox{otherwise}.\end{cases}$
@@
When $r=k$, the preceding bound is equivalent to the lower bound on the access cost for MDS convertible codes proposed by [38, Theorem 9], since the $n,k,d]_{q}$ MDS codes are optimal LRCs with $(k,d)_{a}$-locality.
@@ III Improved Lower Bound on Access Cost
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. For $1\leqslant i\leqslant t$, define
@@
    For each $1\leqslant t\leqslant\tau$, all code symbols indexed by $i\in B_{t}$ are in the same repair set $S_{t}$ of the final code $\mathcal{C}_{F}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant j\leqslant t-1}S_{j}\right)\cap(\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1})\right|<\delta-1\mbox{ when }|B_{t}|<\delta-1.$ (13)
Then,
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta_{i}\leqslant 0,\\ k-\Delta_{i}+|B|,&\mbox{otherwise}.\end{cases}$
@@
where the first inequality holds by $|S_{t}|\leqslant r+\delta-1$ for $t\in[\tau]$, and the second inequality comes from condition (i) that $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for $1\leqslant t_{2}<t_{1}\leqslant\tau$, which implies $|\mathcal{A}_{1}|\leqslant\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$ by $|B|\leqslant(\delta-1)\lfloor\frac{\Delta_{1}}{r+\delta-1}\rfloor$. Note that
@@
Claim 1: We claim that the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\ell}}$ can be recovered by the symbols $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\ell})}$ for $1\leqslant\ell\leqslant\tau$. The claim can be proved by induction as follows. Denote $\overline{\mathcal{U}}\triangleq\mathcal{U}_{2,1}\cup\cdots\cup\mathcal{U}_{t,1}$. If $\ell=1\in\mathcal{A}_{2}$, then $|S_{1}\cap(\mathcal{U}_{1,1}\setminus\mathcal{D}_{1,1})|<\delta-1$ by (13). By (9) and the fact $S_{1}$ is a repair set in $\mathcal{C}_{F}$, we have $S_{1}\subseteq\mathcal{F}_{1}=\mathcal{U}_{1,1}\cup\overline{\mathcal{U}}\cup\mathcal{W}_{1}$. Furthermore, we have
@@
$=$~ $\left|(S_{\ell+1}\setminus B_{\ell+1})\cup\left(\bigcup_{i\in[\ell]}(S_{i}\cap S_{\ell+1})\right)\right|$
$\geqslant$~ $\left|S_{\ell+1}\setminus B_{\ell+1}\right|$
@@
where the first equation holds by $B_{\ell+1}\subseteq S_{\ell+1}\subseteq B\cup N\subseteq M$, the first inequality comes from the condition (i) that $B_{\ell+1}\cap S_{i}=\emptyset$ for $i\in[\ell]$, and the second inequality holds by condition (ii) that $|B_{\ell+1}|\leqslant\delta-1$. Then, the symbols $\bm{v}|_{B_{\ell+1}}$ and $\bm{v}|_{S_{\ell+1}}$ can be recovered by the symbols $\bm{v}|_{M\setminus B_{\ell+1}}$. Therefore, the symbols $\bm{v}|_{B_{1}\cup\cdots\cup B_{\tau}}=\bm{v}|_{B}$ can be recovered by $\bm{v}|_{M\setminus(B_{1}\cup\cdots\cup B_{\tau})}=\bm{v}|_{M\setminus B}$ according to induction.
@@
This implies that $d\leqslant n_{I}-k+1$, which is a contradiction. Thus, when $d>n_{I}-k+1$, we always have $\Delta_{1}\leqslant 0$ and thus $|\mathcal{D}_{1,1}|\geqslant k$, which completes the proof. ∎
The lower bound in Lemma 6 depends critically on the size of the set $B$. Below, we demonstrate that there always exists a set $B$ with large cardinality.
@@
    For each $1\leqslant t\leqslant\tau$, all code symbols indexed by $i\in B_{t}$ are in the same repair set $S_{t}$ of the code $\mathcal{C}$, and $B_{t_{1}}\cap S_{t_{2}}=\emptyset$ for all $1\leqslant t_{2}<t_{1}\leqslant\tau$;
@@
    $\left|\left(S_{t}\setminus\bigcup_{1\leqslant i\leqslant t-1}S_{i}\right)\cap\tilde{V}\right|<\delta-1\mbox{ when }|B_{t}|<\delta-1.$ (15)
@@
where T=∑_1⩽j⩽ℓ-1— ~Vij—¡δ-1 —V_i_j—, and the first inequality holds by the fact that in each round the set $\tilde{V}$ reduces $|V_{i_{j}}|$ for the case $|\tilde{V}_{i_{j}}|<\delta-1$ and at most $|S_{i_{j}}|\leqslant r+\delta-1$ when $|\tilde{V}_{i_{j}}|\geqslant\delta-1$.
@@
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$, and the optimal Hamming distance $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\phi(k,r,\delta)$. Then, the write access cost is
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0,\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$
@@
$|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|\leqslant|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\phi((t-1)k,r,\delta),$
@@
$\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1\leqslant n_{F}-2d-(t-1)k+2-\phi((t-1)k,r,\delta)\triangleq\Delta,$
where $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$ is first defined in Lemma 6. Thus, $\Delta_{i}\leqslant\Delta$ for every $i\in[t]$. According to Lemma 7, let $\tilde{V}=\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}$ then $|\tilde{V}|=\Delta_{i}+d-1$ by $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{D}_{i,1}|-d+1$. Then, we can find a $B$ satisfying conditions (i), (ii) and $|B|=(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$. This is to say that
@@
$|\mathcal{D}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta\leqslant 0,\\ k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,&\mbox{otherwise},\end{cases}$
@@
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0,\\ t\left(k-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right),&\mbox{otherwise}.\end{cases}$ (23)
@@
Remark 2:
Corollary 1 extends the lower bound in Theorem 1 [42, Theorem IV.4] which is only associated with $r$, to one related to $r$ and $\delta$. For the case $\delta=2$, the new lower bound is exactly the same as the one in Theorem 1. In the case $\delta>2$, the new lower bound on the write access cost clearly outperforms the one in Theorem 1. The expressions $\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor,\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$ are not easily reducible and LRCs achieving the minimum Hamming distance bound in Lemma 2 usually require some parameter constraints [50, 51], e.g., $r+\delta-1\mid n$, or $n=w(r+\delta-1)+m,k=ur+v$ such that $0<v<r,v+\delta-1\leqslant m<r+\delta-1$, etc. Therefore, we investigate the following specific scenarios which show the new lower bound on the read access cost surpasses the one in Theorem 1, i.e., $\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\leqslant\left\lceil\frac{r\overline{\Delta}}{r+1}\right\rceil$:
  1. 1.
    When $\Delta\mid(r+\delta-1)$, then
  2. 2.
    Let $\bm{\mathcal{C}}$ be with parameters $(n_{I},kr,*,*;n_{F},tkr,r,\delta)_{a}$ and the final code $\mathcal{C}_{F}$ be an optimal LRC satisfying $n_{F}=(tk+m)(r+\delta-1)$, where $m$ is a positive integer and $tk-m\geqslant 2$, then
The reader may refer to Figure 1 as an example.
Figure 1: Compare the lower bounds on read access cost of Theorem 1  [42, Theorem IV.4] and Corollary 1. We consider only the case $\Delta>0$ and $d\leqslant n_{I}-k+1$. Set the $t=3$ initial codes with parameters $n_{I}=80,k=40]_{q}$, and the final code with parameters $tk/\alpha,tk,d=15]_{q}$ and with $(r=10,\delta)_{a}$-locality where $\alpha$ is the code rate (the X-axis in the figure) of the final code. Note that the lower bound of Theorem 1 is independent of the $\delta$, therefore, the conditions for $\delta=3,4,5$ are the same blue circle in the figure.
Remark 3:
In the following, we provide lower bounds for the access cost under two classes of parameters. These bounds correspond precisely to the parameter sets constructed in subsequent sections.
@@
Suppose $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ corresponding to Lemma 2. Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $((\ell/t+g-\ell)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$, where $t\mid\ell$, $\ell/t>g-\ell\geqslant h\geqslant 0$ and $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d=d_{\mathrm{opt}}^{(q)}$. Then the write access cost is at least $h(r+\delta-1)$ and the read access cost is at least thrthr.
Proof.
Based on Corollary 1, the write access cost $\rho_{w}$ satisfies
$=t(\delta+h(r+\delta-1)+(t-1)r\ell/t-1+((t-1)\ell/t-1)(\delta-1))$
$\quad-(t-1)(\ell+h)(r+\delta-1)$
$=h(r+\delta-1),$
and
$=(\ell+h)(r+\delta-1)-2(\delta+h(r+\delta-1))$
$\quad-(t-1)r\ell/t+2-((t-1)\ell/t-1)(\delta-1)$
$=(\ell/t-h-1)(r+\delta-1)+r$
$>0,$
where the last inequality holds by $\ell/t>g-\ell\geqslant h$. Therefore, the read access cost $\rho_{r}$ satisfies
$\rho_{r}$ $\geqslant t\left(r\ell/t-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right)$
$=t(r\ell/t+h(r+\delta-1)-(r+\delta-1)\ell/t+\delta-1+(\delta-1)(\ell/t-h-1))$
$=thr.$
Corollary 3:
Suppose $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$ corresponding to Lemma 3. Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $((\ell/t+h)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$, where $\delta\geqslant 2,t\mid\ell$ and $\ell/t>h\geqslant 0$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d=d_{\mathrm{opt}}^{(q)}$. Then the write access cost is at least $h(r+\delta-1)$ and the read access cost is at least thrthr.
@@
Corollaries 2, 3 show the same lower bounds on access cost, although the value of $d_{\mathrm{opt}}^{(q)}$ is not the same. This inspires us to use the same method when constructing the access-optimal LRCCs with $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1)$ in Section V-B or $d_{\mathrm{opt}}^{(q)}(n,k,r,\delta)=n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1)$ in Section V-A.